### v2

# X-SG2S: Safe and Generalizable Gaussian Splatting with X-dimensional Watermarks

Abstract.
3D Gaussian Splatting (3DGS) has been widely used in 3D reconstruction and 3D generation. Training to get a 3DGS scene often takes a lot of time and resources and even valuable inspiration. The increasing amount of 3DGS digital asset have brought great challenges to the copyright protection. However, it still lacks profound exploration targeted at 3DGS. In this paper, we propose a new framework X-SG2S which can simultaneously inject 1 to 3D watermarks while keeping the original 3DGS scene almost unchanged. Generally, we have a X-SG2S injector for adding multi-modal messages simultaneously and an extractor for extract them. Specifically, we first split the watermarks into message patches in a fixed manner. A self-adaption gate is used to pick out suitable location for watermarking. Then use a XD(multi-dimension)-injection heads to add multi-modal messages into sorted 3DGS points. A learnable gate can recognize the location with extra messages and XD-extraction heads can restore hidden messages from the location recommended by the learnable gate. Extensive experiments demonstrated that the proposed X-SG2S can effectively conceal multi modal messages without changing pretrained 3DGS pipeline or the original form of 3DGS parameters. To more comprehensively evaluate the generalizability and practicality of our model, we assessed it from several aspects: the ability to maintain watermark and original 3DGS consistency, robustness against model degradation, and the model’s ability to make correct judgments. The experimental results demonstrate that our model performs exceptionally well in these areas. Meanwhile, with simple and efficient model structure and high practicality, X-SG2S still shows good performance in hiding and extracting multi-modal inner structured or unstructured messages. X-SG2S is the first to unify 1 to 3D watermarking model for 3DGS and the first framework to add multi-modal watermarks simultaneous in one 3DGS which pave the wave for later researches.
3DGS, Multi Modal Watermarks, Copyright Preservation
ccs: Security and privacy Social aspects of security and privacy
1. Introduction
Figure 1. Application scenario of the proposed. Scenario 1: after training a 3DGS pipeline by a trainer, an X-SG2S injector should be combined for joint use. In this way, when other users employ the trainer’s model, they can obtain a 3DGS result with some watermarks. By providing the 3DGS file to X-SG2S’s extractor, the trainer can determine whether a user has generated the result by using the trainer’s model. Scenario 2: if a user has specific requirements to add extra information to the 3DGS result for a particular purpose, they can directly use X-SG2S to achieve this goal. For example, you can add a text with “ACMMM”, an image of the ICML logo and a small 3DGS object at the same time by using X-SG2S’s injector and extract them by using X-SG2S’s extractor.
The applications of 3D reconstruction and 3D generation technology are extremely extensive. It has already made significant contributions in fields such as healthcare, architecture and engineering, cultural heritage preservation, virtual reality and augmented reality, as well as game and film production. The existing work mainly studies how to build 3DGS models to generate 3D scenes or reconstruct 3D scenes.
However, the increasing emergence of 3D Gaussian Splatting (3DGS) models may lead to issues of information leakage and infringement. At the same time, in the near future, it is likely that people will need to add more, and even multimodal, information or watermarks to a reconstructed or generated scene which poses a challenge to existing methods. Meanwhile, using different methods to add watermarks makes the watermarking adding mechanism complicated and chaotic.
Currently, only a small portion of research has addressed these problems. The new studies GS-Hider(zhang2024gs, ) and Splat in Splat(guo2024splats, ) restore image or 3D scene into another 3D scene. Considering the security issues of generative 3D, GaussianStego(li2024gaussianstego, ) explores how to embed image watermarks into diffusion models. GaussianMarker(huang2024gaussianmarker, ) embed 1D watermark into 3DGS. However, these models all have their limitations. For example, GS-Hider modifies the parameters of GS, which may increase the size of the 3D model parameters and no longer make it universal. Splat in Splat and GS-Hider can add watermarks only when they have the whole dataset of two scene. And they cannot add watermarks when user only have 3DGS model and watermarks message. GaussianMarker can only inject little binary message into 3DGS. The approach of GaussianStego rooting dino into diffusion can not well demonstrate its effect in feed forward 3D generative models. At the same time, adding extra watermarks in original 3DGS pipelines for protecting copyright is costly and complicate in design or training. We can see that how to effectively add additional watermarks without changing the GS parameters or model structure and let the method be general will be a great challenge. Meanwhile, embedding multimodal watermarks into a 3DGS simultaneously is even more of an unexplored area.
To solve the above challenges, we propose an effective, simple and flexible watermarking framework X-SG2S. It aims to embed 3D objects or images or binary messages simultaneously into the original 3DGS scene, and accurately extract the hidden message via just an ultra simple way. Specifically, we use patch the raw data and create redundancy which can let the watermark extracting process more robust. We innovatively proposed a model feature sparse DCAE for the restoration image when the some features are loss. Then, we use a self-adaption gate which can learn the adding position through the interaction between the additional watermarks and the 3DGS scenes. A XD-injection heads are used to add different modal messages respectively. Then a learnable gate is used to recognize the locations where we have added the messages. Finally, XD-extraction heads are used to restore the message added. The training for X-SG2S does not need to fine tune or redesign the pipeline. This property makes this model suitable not only for the feedforward model, but also for the model like diffusion which needs multiple iterations. And if you do not need to add additional information, you can use the original model in the fastest time without further fine-tuning. However, it is also possible to fine tune with the pipeline. Extensive experiments demonstrated that the proposed X-SG2S can effectively conceal multimodal messages without compromising rendering quality and possesses exceptional fidelity, robustness, veracity and flexibility. In a nutshell, the contributions and advantages of our X-SG2S can be summarized as follows.
  • We have pioneered the first method that can simultaneously embed watermarks of 1 to 3 dimensions into one 3DGS scene, which paves the way for future research.
  • We have provided an effective approach for GS clouds to store additional structured or unstructured data.
  • We create a model that can restore the whole image when some features are missing.
  • Our model, in an extremely concise and efficient manner, makes the process of adding watermarks to 3DGS both universal and secure.
  • We conducted extensive experiments to demonstrate the efficiency. The experimental results were SOTA, which established the baseline method for later research.
2. Related Works
2.1. Generalizable GS
3D Gaussian Splatting (3DGS) (kerbl20233d, ) has emerged as a powerful method for reconstructing and representing 3D scenes using millions of 3D Gaussians. There has been a significant amount of work that has achieved good results in the reconstruction of small objects (szymanowicz2024splatter, ) (boss2024sf3d, ) and scenes (chen2025mvsplat, ) (charatan2024pixelsplat, ). Additionally, some studies have utilized diffusion models to realize text-to-3D generation (li2024instant3d, ), which also achieves promising outcomes.
2.2. 3D Steganography
Steganography has been evolving over the decades. Some research has made 3D steganography achieve good results in explicit geometry like meshes and point clouds (ohbuchi2002frequency, ) (zhu2024rethinking, ) (ferreira2020robust, ) or implicit geometer like nerf (li2023steganerf, )(luo2023copyrnerf, ). 3DGS is a new technology to represent explicit geometry. However, few works have been researched in steganography for 3DGS. GS-Hider (zhang2024gs, ) design a coupled secured feature attribute to replace the original 3DGS’s spherical harmonics coefficients and then use a scene decoder and a message decoder to disentangle the original RGB scene and the hidden message. Splat in Splat (guo2024splats, ) first train two scene by using the same location and shape of the cover scene and different color and opacity. Then they inject the secret scene into the cover one by an unlearnable method. And use an AE to map the transparency variation of each point. This method can work only when you have the entire multi view datasets. And for every scene, you should learn a new AE which is very fussy. GaussianStego(li2024gaussianstego, ) use dino to add image watermarks and use U-Net to extract them via multi rendered views. But there is not research about watermarking 1 to 3D messages to single 3DGS scene simultaneous while keeping GS parameters or model structure unchanged. GaussianMarker(huang2024gaussianmarker, ) can only embed 48 bit watermarks. The added information might be insufficient in specific contexts.
3. Safe Generalizable GS with X-dimensional watermarks
3.1. Preliminaries of Gaussian Splatting
Gaussian Splatting is a rasterization technique for 3D scene reconstruction and rendering. It works by using 3D Gaussian functions to represent points in a scene and projecting these Gaussian functions onto a 2D image plane for rendering. For every point, it has many different parameters to represent the location, shape, size, color and opacity.
$\exp\left(-\frac{1}{2}(\boldsymbol{x}-\boldsymbol{\mu}_{i})^{\top}\Sigma_{i}^{-1}(\boldsymbol{x}-\boldsymbol{\mu}_{i})\right)$
This function shows the location, shape and size of a GS point where $0<i<N$, $\mu_{i}\in R^{3}$ is the mean or center, $\Sigma_{i}\in R^{3\times 3}$ is its covariance specifying its shape and size. The covariance can be decomposed into scaling factor $s\in R^{3}$ and rotation factor $q\in R^{4}$. Each Gaussian has also an opacity $\alpha_{i}\in[0,1]$ and a view-dependent color $c_{i}(v)\in R^{3\times(k+1)^{2}}$ where k is the higher level of spherical harmonics(SH). The set of GS points G = {($\mu_{i}$ $\Sigma_{i}$ $\alpha_{i}$ $c_{i}$), i =1,…,N} can be used to represent a scene or a small object. Through rasterization, it can be rendered to an image where you only need to know the projective transformation P, viewing transformation W, and Jacobian of the affine approximation of P, J, respectively.
$\hat{\mu_{i}}=PW\mu_{i}$
$\hat{\Sigma_{i}}=JW\Sigma_{i}J^{T}W^{T}$
$\sigma_{i}=\alpha_{i}\mathrm{e}^{-\frac{1}{2}(\mathbf{p}-\hat{\boldsymbol{\mu_{i}}})^{\top}\hat{\boldsymbol{\Sigma_{i}}}^{-1}(\mathbf{p}-\hat{\boldsymbol{\mu_{i}}})}$
$\mathbf{C}[\mathbf{p}]=\sum_{i=1}^{N}\boldsymbol{c}_{i}\sigma_{i}\prod_{j=1}^{i-1}(1-\sigma_{j})$
where p represents the pixel.
3.2. Task Settings
3DGS steganography: We hope that for any 3dgs scene, we can add additional information without damaging the original 3d structure and gs parameters.
Copyright Protection: By adding fixed watermark into generated scenes, the designers can easily use X-SG2S to detect whether the GS art works are used by others without permission.
Multi-modal Message Adding: This model can not only add information of single modality, but also add information of multiple modalities without producing confusion. This can greatly bring convenience to the users. In theory, this model can add sequential structure data or non-sequential structure data, and accurately extract them when the GS point set is well preserved.
Generalizable Modules: This model does not require changing the structure of the GS generated model, but rather adds watermarks or additional information to the scene in the form of additional modules.
3.3. Deep Thinking About Watermarking 3DGS
Gaussian sphere is a special kind of point set. You can simply see it as a kind of point clouds with many parameters. First of all, we think about why it is difficult for us to add information to GS point clouds.
  • Point set is an unordered set which means you may treat the point set as a whole to add information (e.g. add information to the frequency domain of the entire point cloud or add the information to the entire point cloud). But this can make the waste of storage space when the amount of information is small or may make large damage to the original model.
  • If you just use a part of points to add information, you have difficulty knowing the location and order ,especially, when you add more than one watermarks.
  • Each point has a limited capacity to store information. How to correctly use small containers to store watermark with large amount of information is also a challenge.
After deep thinking, we can sort out such information adding ideas:
  • Every point in the point set have a few of storage space. But when all the storage units are added up, the storage capacity is very large. If we really want to fully use all of the storage space, we should smash the information into patches in an order that we can restore them easily and add them respectively for every points. If we add too much information to one of the point, it may cause loss.
  • If we want to know how to extract the information from the point set, we should firstly know which of them are added. This means you would better know the location of each point and the restoration order.
3.4. Overview
Figure 2. This figure shows the structure of self-adaption gate.
The major target of X-SG2S is to design a general framework for multi-modal message adding model which can inject and extract 1D,2D and 3D massages with one model while keeping the original Gaussian scene as unchanged as possible.
As depicted in Figure 5, X-SG2S contains mainly four components: 1) an adaptive self-adaption gate for adaptively selecting GS points to use. 2) a multi-head MLP-based message injector for integrating information into SH parameters. 3) a learnable selection gate for recognizing the location of injected GS points. 4) a multi-head MLP-based message extractor for restoring message from the injected GS points. X-SG2S injector is made up of 1) and 2), extractor is made up of 3) and 4).
The training process and the inference process are slightly difference. The following part will illustrate them in detailed.
3.5. Patch the raw data and increase the number of patches
We need to patch the raw data because they usually have large size which can not be injected to single point. However, if we just inject the raw data patches, it will let the watermarks vulnerable to attack. If some of the GS points are hostile removed, we have no choice to restore all the data patches which means the watermarks will be destroyed. We need to find out a method to create patches from raw data, which can restore the whole watermark when the scene is attacked.
As shown in the Figure 4, for 1D and 3D data, we just duplicate the raw date many times because they are already patches. By repeating them, we can restore the watermarks by any subset of these data without difficulties. But for 2D message, things may be different. If we directly split the image into patches, we will get a large number of patches. At the same time, we cannot restore the whole image by any of the subset. So, we introduce a new method, feature sparse DCAE, which can restore the whole image by using subset of feature patches.
In the Figure 3, we show our feature sparse DCAE. Thanks for powerful ability of DCAE(chen2024deep, ), we can simply encode one 256 revolution image into a feature of shape $4\times 4\times 128$. Then, we flatten the feature and see them as a sequence. We use 4 up sampling convolution layers to create redundant context of this sequence and add fixed position encoding. In order to simulate the hostile attack, we ramdomly remove some of the patches. But this time, because we have a sequence with context, we have the ability to restore the whole sequence by any of the subset. Specifically, we using a MLP to predict the fixed position encoding, then use the predict encoding and the complete encoding to do nearest neighbor algorithm which find out where should the patches really located. We let all the patches be placed in the correct position. Then we use down sampling convolution layers and refinement MLP to restore origin feature sequence. Finally, we rearrange them and decode the watermark.
Figure 3. This figure shows the structure of feature sparse DCAE.
Figure 4. This figure shows we how to redundantize and patch the raw data to let the watermarking more robust.
3.6. Self-Adaption Gate
This component is to find out the points which are suitable for injecting the messages while ensuring the original scenario is not disturbed so much. A simple idea is that let a gate to learn a score to measure the locations. The gate will give every point a comprehensive score. A higher score indicates that the location is more suitable for adding information.
The comprehensive score is calculated from two scores:“self-score” and “cross-score”. “Self-score” is used to measure which points can significantly impact the 3DGS scenario and which are trivial. But only use “self-score” is not enough. We need to know where are more suitable for message to inject and extract. “Cross-score” is used to manage this problem.
After we get two scores, we simply multiply two scores together to get the final score. By using top-k, we can set the recommended k positions to 1 and others to 0 to get the selection mask. The structure is shown in Figure 2.
3.6.1. message embedding
Different modalities belong to different vector spaces. If we want to calculate the interactions results uniformly and efficiently, we need to map them into a unified space first. Here we simply use three different linear layer to map 1,2,3D and GS cloud message to a general space.
3.6.2. induced set attention to produce self-score
After we get embedded message, we can measure the interact scores. Attention mechanism is a very efficient method for interact different information. We use transformer-based model to calculate the “self-score”. But we find that if we simply use scale dot product attention, the memory will overflow because the number of GS points are large. However, to calculate “self-score” is merely want to measure the importance of individual information in the overall information instead of learning detailed information between individuals. So set attention(lee2019set, ) is used to overcome the memory problem. At the same time, set transformer does not need for the data to be sequential. This is very suitable for point sets, because point sets are unordered.
Induced set attention first design a set of anchor vectors $I\in R^{m\times d}$, where d means the inner dimension of the vector, m means the number of anchor vectors. Then it uses the input $X\in R^{n\times d}$ and I to do multi-head transformer to get the hidden message $H\in R^{m\times d}$ of the point set. Then use H and X to do multi-head transformer again to get the importance of every points in overall information $O_{s}\in R^{n\times d}$. Finally, use $O_{s}$ to calculate the $self\_score\in R^{n\times 1}$ by using a linear layer. Formally,
$H=MA(I,X)$
$O_{s}=MA(X,H)$
$self\_score=sigmoid(linear(O_{s}))$
In our task, we simply set the number of heads to 1 and the number of anchor vectors to 1 so that the time complexity goes from $O(N^{2})$ to $O(N)$.
3.6.3. efficient attention to produce cross-score
Because “Cross-score” needs to interact the whole GS points with embedded message patches. However, if the amount of message patches is large, the attention may also face the problem of overflowing. Thus, we use efficient transformer (shen2021efficient, ) to approximation the common transformer in order to reduce the memory and preserve the accuracy of calculating.
Because this is a cross attention mechanism, we let $X$ to become $Q$, and the message patches $Y\in R^{l\times d}$ to become $K$,$V$. It is easy to get that $l$ is usually less then $n$. So, it is efficient to let $KV$ interact first. Formally,
$O_{c}=\sigma_{row}(Q)(\sigma_{col}(K)^{T}V)$
Where $\sigma$ means the softmax function. The same as above, use $O_{c}$ to calculate the $cross\_score\in R^{n\times 1}$ by using a linear layer.
$cross\_score=sigmoid(linear(O_{c}))$
3.6.4. Location masks
After getting two scores, we simply multiply two scores to get a final score. For every modality, it has a final score. This final score can comprehensively evaluate whether a point is suitable to add message. We use top-k to select the most suitable locations for adding, where k is the number of patches. If we want to generate the next modality’s location mask, we should use the previous masks to avoid selecting the duplicate locations.
3.7. Learnable Selection Gate
For above we can know that the self-adaption gate can produce a mask. Here we see the masks as the ground truth. The learnable selection gate see ordered point set with added messages as the input to find out the locations where the information added. The gate is a four layers MLP to predict the true location.
3.8. Importance of SH
According to the Splat in Splat(guo2024splats, ), different order of SH has different importance. Higher-order SH contains only a small amount of High-frequency information and has less contribution to the scene. And we also do the experiments to prove this concept. So, in our machine, we choose to add watermark patches to order 3 and 4 of the SH.
3.9. XD-Injection/Extraction Heads
These two are used for injecting messages and extracting messages. For each modal, we use a separate MLP (qi2017pointnet, ). Usually, 3 to 5 layers are enough for the task. We take out the GS points of target location recommended by self-adaption gate and just use the spherical harmonic parameters of three channels. We concatenate the spherical harmonic parameters with the message patches we want to added. We treat the concatenated data as input to multi-head injector which produce the spherical harmonic parameters with extra information.
And for extractor, it take the SH parameters of injected points found out by learnable selection gate as input and extract the messages from them. The extracted messages will be optimal by the prediction loss as well.
3.10. Training Process
Figure 5. This figure shows the structure and training process of X-SG2S. Orange lines mean there exists gradient flow. Blue lines mean the gradient are truncated. Purple lines mean the GS cloud is sorted by sorting method. Black lines mean the losses used to train.
This model can be trained directly by GS cloud files or a pretrained 3DGS pipeline. In Figure 5, we show this process.
  1. (1)
    Get the GS cloud inferred from a pipeline or read from a cloud file.
  2. (2)
    Preprocess the raw data into patches.
  3. (3)
    Give self-adaption gate the message patches and the GS cloud to recommend the locations.
  4. (4)
    Use XD-injection heads to add message.
  5. (5)
    Replace the original SH parameters by the injected SH parameters so that we get a GS point set with extra messages.
  6. (6)
    Truncate the gradient flow of the GS point set with extra messages and take it as input to the learnable gate. It will output the location mask. This part is off-line to ensure proper flow of gradients. If we directly choose the predicted location to extract messages, the message extractor may choose some GS points without messages. This may cause the error of gradient flow.
  7. (7)
    Use the truth location to choose the injected GS points and extract messages by XD-extraction heads.
3.11. Training Loss
To finish the task, we should design 1) message patches loss, 2) location mask loss, 3) SH parameters loss, 4) other losses.
To ensure the feature patches can be restore exactly, we should use message patches loss to ensure the extracted message is as same as possible to the message injected. For 1D message, we use cross entropy loss; for 2D and 3D message, we use MSE loss.
To ensure the GS parameters do not change much before and after adding information, we use a SH parameter MSE loss to ensure this.
To ensure the learnable selection gate can precisely recognize the location where we inject the message, we use a mask cross entropy loss. The weight for the mask cross-entropy loss does not need to be set as its training is offline. We default this weight to 1.
Other loss like 2D view loss and lpips loss can be choosed to use when we use pretrained 3DGS pipeline to train X-SG2S. They may let X-SG2S more efficient to preserve the SH parameters.
Thus, the form of loss function is like:
[…]
4. Experiments
4.1. Datasets
We use a pretrained model MVSplat for real time inference. We use large-scale ACID datasets to let MVSplat(chen2025mvsplat, ) generate the 3DGS scenes. For 1D dataset, we randomly generate a sequence of binary message. For each bit, it has the same probability to be 1 or 0. For 2D dataset, we use Logo-2K(wang2020logo, ) dataset. For 3D dataset, we first download a subset of objaverse(deitke2023objaverse, ), and use Gamba(shen2024gamba, ) to generate the 3DGS objects. They are then saved in .ply format as a 3D dataset.
Specifically, ACID contains nature scenes captured by aerial drones, which are split into 11,075 training scenes and 1,972 testing scenes. After reconstructing Logo-2K dataset, it has 110313 training logos and 28415 testing logos. 3D dataset made by objaverse and Gamba contains 353 training files and 94 testing files. To ensure the number of datasets is aligned, we reuse the GS cloud files while training.
4.2. Metrics
For original GS scenes and GS objects, we use pixel-level PSNR, patch-level SSIM, and feature-level LPIPS to measure the completeness of the original scenes and the quality of restoration of the GS objects. For a fair comparison, all the scenes are rendered to 256$\times$256 resolutions while all the objects are rendered to 512$\times$512 resolutions. For 1D data, we use precision rate to measure the extraction accuracy of binary data. For 2D data, we use pixel-level PSNR to quantization the restoration of the 2D images.
4.3. Implementation Details
X-SG2S is implemented with PyTorch, along with an off-the-shelf 3DGS render implemented in CUDA. All models are trained on single RTX4090 with the Adam optimizer.
Its training involves multimodality, so there are many hyperparameters in the loss function that need to be set. We default to using optional loss functions unless in loss function ablation experiments. The optional loss functions include 2D rendered image MSE loss and LPIPS loss, with weights of 1 and 0.05, respectively. And other hyperparameter are $\gamma=0.2$, $\phi=0.005$, $\theta=0.8$, $\delta=1.5$.
We choose binary message of 48bit×$\times$1024 as 1D watermark and 3DGS object with 10000 points as 3D watermark. For 2D watermark, we use feature sparse DCAE to get 16$\times$2048 message patches. And the number of higher order SH parameters is 48 for each GS point.
4.4. Main Results
4.4.1. ablation experiments
Figure 6. Interaction Blocks: a) “Multiplication”: it multiplies the self-score and the cross-score to generate final-score. b) “Addition”: it uses efficient cross attention to generate cross-feature and set self attention to generate self-feature. Then it concatenates them and calculates final-score. c) “Concatenation”: it concatenates cross-feature and self-feature together and calculates final-score. d) “Normal”: it just uses 3DGS original points as input and passes them through n layers of MLP gate. e) “Random”: it randomly selects k locations to add extra message. It can not be learned.
In Figure 6, we design five different structures of interaction block and conduct ablation experiments respectively. We jointly train X-SG2S with multimodal watermarks. In Table 1, we show that the method of a) is the best.
Method Multiplication Addition Concatenation Normal Random
PSNR(org)$\uparrow$ 27.666 25.812 25.261 25.951 25.614
SSIM(org)$\uparrow$ 0.868 0.839 0.833 0.845 0.759
LPIPS(org)$\downarrow$ 0.135 0.214 0.208 0.203 0.246
$\uparrow$ 1.0 1.0 1.0 1.0 1.0
PSNR(image)$\uparrow$ 30.537 30.421 29.991 30.174 30.552
PSNR(3D obj)$\uparrow$ 21.140 21.794 21.201 21.141 21.231
SSIM(3D obj)$\uparrow$ 0.843 0.846 0.855 0.842 0.848
LPIPS(3D obj)$\downarrow$ 0.166 0.159 0.163 0.157 0.158
Table 1. Multimodal joint training with different structures of interaction block.
Method 1D-SG2S 2D-SG2S 3D-SG2S Origin
PSNR(org)$\uparrow$ 27.837 27.811 27.759 28.25
SSIM(org)$\uparrow$ 0.871 0.866 0.870 0.843
LPIPS(org)$\downarrow$ 0.125 0.131 0.133 0.144
$\uparrow$ 1.0 / / /
PSNR(image)$\uparrow$ / 30.537 / /
PSNR(3D obj)$\uparrow$ / / 21.144 28.807
SSIM(3D obj)$\uparrow$ / / 0.843 0.942
LPIPS(3D obj)$\downarrow$ / / 0.165 0.067
Table 2. Single-modality watermark injection and extraction.
4.4.2. single modality injection test
In Table 2, we test our X-SG2S on adding sigle watermark. We can see that X-SG2S can well adapt the task of 1 to 3D watermarking. We can see that the performance of watermark extraction are almost identical. This demonstrates the effectiveness of our method.
4.4.3. The Exploration of Loss Function
Method without optional losses with optional losses
PSNR(org)$\uparrow$ 25.241 27.666
SSIM(org)$\uparrow$ 0.842 0.868
LPIPS(org)$\downarrow$ 0.172 0.135
$\uparrow$ 1.0 1.0
PSNR(image)$\uparrow$ 29.132 30.537
PSNR(3D obj)$\uparrow$ 20.616 21.140
SSIM(3D obj)$\uparrow$ 0.837 0.843
LPIPS(3D obj)$\downarrow$ 0.166 0.166
Table 3. Comparison with and without optional loss functions.
Now, we want to know if X-SG2S can be trained without optional losses which means it can trained by using just 3DGS files. Here we set the weights of optional losses to 0 and set the $\gamma=2$. From Table 3, we can see that this way is feasible! However, X-SG2S trained without optional loss functions performs worse in both watermark embedding effectiveness and preserving the integrity of the original scene compared to those trained with optional loss functions. Therefore, we recommend using optional loss functions to enhance the effectiveness of the training.
4.5. Fidelity
According to Figure 7, we show the importance of SH. The left column shows X-SG2S injects the watermark to all the SH while the right shows the watermarks are only injected to the higher order SH. It seems that only inject watermarks to higher order SH can achieve good results. From the visual perspective, the invisible watermark cannot be detected by the naked eye which means we protect the original scene well and hide the watermarks well. Meanwhile,we have drawn the residual map of the invisible watermark in Figure 8. It can be observed that our method can effectively hide the watermark without detection. Figure 9 shows that our X-SG2S can inject and extract watermarks well.
Figure 7. This image illustrates the different effects of watermarking all SH coefficients versus watermarking only the high-order SH coefficients.
Figure 8. This image demonstrates the effectiveness of our approach. The first column shows the original scene images rendered by MVSplat. The second column displays the images after watermark embedding. The third column presents the 10×residuals between the two images. It can be seen that our model is capable of effectively concealing the watermark.
Figure 9. This image demonstrates the watermarks. The first column shows the original watermarks provide by Logo-2K. The second column displays the watermarks extracted by X-SG2S. The third column presents the GS objects extracted by X-SG2S. The final column displays the GS objects provided by Gamba.
4.6. Robustness
To evaluate the robustness of X-SG2S, we have subjected the Gaussians to degradation using random pruning methods. Quantitative metrics are shown in Tab 4. The results indicate that our method effectively withstands the degradation process.
Pruning Rate PSNR(org) SSIM(org) LPIPS(org) precision PSNR(image) PSNR(3D obj) SSIM(3D obj) LPIPS(3D obj)
5% 26.799 0.821 0.246 1.0 30.511 20.568 0.843 0.169
10% 25.784 0.775 0.294 1.0 29.847 20.572 0.842 0.170
15% 24.661 0.728 0.333 1.0 29.931 20.520 0.842 0.171
20% 23.485 0.682 0.366 1.0 29.620 20.484 0.842 0.173
25% 22.273 0.635 0.398 1.0 29.571 20.433 0.841 0.174
Table 4. Robustness testing by random pruning GS points.
PSNR(org) SSIM(org) LPIPS(org) precision PSNR(image) PSNR(3D obj) SSIM(3D obj) LPIPS(3D obj)
GS without wm 28.25 0.843 0.144 0.395 7.719 11.934 0.835 0.319
Table 5. To verify whether the detection model can distinguish if the original model contains a watermark.
4.7. Precise Identification
In the real case, we cannot indiscriminately detect watermarks. That is to say, we cannot falsely identify a watermark in a model that does not have one. To illustrate this, we randomly select some GS scenes without watermarks and observe that our model fails to detect any watermark. Table 5 demonstrates that our model can effectively reduce false positives without the need for an additional discriminator.
5. Conclusion
In this work, we propose a multi modal watermarking framework for 3DGS , X-SG2S, designed for efficient adding 1,2,3D extra messages to a GS scene while keeping the original scene almost unchanged. X-SG2S successfully adds binary, graph, 3DGS object messages to a 3DGS scene simultaneously or individually. Additionally, X-SG2S does not need to fine tune pretrain model (but train with pretrained models is workable) and it can even trained on only 3DGS files which depletes training time and space. Besides, X-SG2S does not change the form of GS parameters so that it is suitable for most of 3DGS scenes or objects.